what is the x-intercept of the line 5x-3y=-15
Answer by Alan3354(69443)

what is the x-intercept of the line 5x-3y=-15
It's the point where the line crosses the x-axis. That means y = 0.
--------
5x-3y = -15
5x-3*0 = -15
5x = -15
x = -3
The point is (-3,0)